OpenStudy (anonymous):

If a = 2 and b = -3, evaluate (a2 - b2)and^3

OpenStudy (anonymous):

you mean: \[(a ^{2}-b ^{2})^{3}\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

(4-9)^3=-5^3=-125

OpenStudy (anonymous):

thanks